OpenStudy (anonymous):

solve the system of equations by substitution... f-3g=9 5f+g=-3

OpenStudy (anonymous):

do you mean f-3g=9 ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Well, \[g=-5f-3\], according to your information. So I substitute that into \[f-3g=9\]. That means \[f-3(-5f-3)=9\] or \[f+15f+9=9\]. That means \[15f=0\]. So \[f=0\]. Now I substitute 0 in for f: \[g=-5(0)-3\]. So \[g=3\] and \[f=0\] :).

OpenStudy (anonymous):

g=-3*